# A framework for the valuation of insurance liabilities by production cost

### v1

@@ 4.1 Non-negative production strategies
We consider insurance liabilities $\mathcal{L}$ from insurance contracts written by an insurance company with outgoing contractually specified cash flows $X_{t}^{\mathcal{L}}\geq 0$ for claims and benefits and other (e.g. administrative) costs and actual incoming cash flows $\tilde{Z}_{t}^{\mathcal{L}}\geq 0$ for premiums, all at dates $t\in D[0,T]$, where the liabilities are extinguished at date $T$, with no remaining cash flows or value. Our objective is to value the insurance liabilities $\mathcal{L}$ by the cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ of strategies $\phi$ that produce them, using capital $\mathcal{C}$ and in general also illiquid assets $\psi$. In the central Definition 4.4 below of non-negative production strategies, the liabilities are produced by a strategy $\phi$ by trading the tradables (in $\phi$) and paying the liability cash flows $X_{t}^{\mathcal{L}}$ at date $t\in D[0,T]$ by selling tradables or paying out cash flows that are incoming at date $t$ from insurance contracts ($\tilde{Z}_{t}^{\mathcal{L}}$), tradables ($\tilde{Z}_{t}^{\phi}$) or illiquid assets ($\tilde{Z}_{t}^{\psi}$). The value $\bar{v}_{i}^{\psi}(\mathcal{L})$ of the liabilities $\mathcal{L}$ given the illiquid assets $\psi$ is then defined as the minimum (or infimum) of the production cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ over strategies $\phi$ and capital $\mathcal{C}$ with fixed illiquid assets $\psi$ (Section 4.4). Production and valuation are for all insurance liabilities of a company together, assuming that no additional insurance liabilities are taken on in the time period $[0,T]$. We do not explicitly model policyholder options to terminate the contracts and receive a ”liquidation value” cash flow (similar to e.g. Definition 3.3) and instead assume that the exercise of such options is taken into account in the cash flows. The extension to not necessarily non-negative strategies ϕitalic-$\phi\in\mathcal{R}^{\prime}$ requires additional assumptions and is provided in Section 4.5.
Insurance liabilities are produced by an insurance company subject to solvency regulation and insolvency laws. Regulatory solvency capital requirements with a one-year view in solvency systems such as Solvency II ([3], [6]), the IAIS International Capital Standard (ICS), and the SST ([9]), e.g. under Solvency II given by a $99.5\%$ value-at-risk, can be seen to correspond to the following requirement set out below in Definition 4.2: the insurance liabilities do not need to be produced always but “in sufficiently many cases” according to a “fulfillment condition” that applies for each successive one-year period ($i$ to $i+1$). E.g. under Solvency II roughly with at least $99.5\%$ probability over each one-year period. The insolvency laws take over when failure occurs, and we define failure such that there is no failure precisely in the “sufficiently many cases” in which the liabilities are fulfilled. This is expressed by the condition that the value of the assets is at least as large as the value of the liabilities, $A^{\prime}_{i}\geq L_{i}$ according to Definition 4.1 below. Our definition of failure is based on specific assumptions about insolvency laws, in particular the relevant valuation standard. These and the further assumptions made here are discussed further below; they may not fully conform with applicable insolvency laws in specific jurisdictions. In Section 4.2, the definition of non-negative production strategies from Definition 4.4 below is extended under specific assumptions from the “sufficiently many cases” to also cover failure and thus all cases.
@@
Let $\phi$ be a non-negative production strategy (Definition 4.4) for the liabilities $\mathcal{L}$ using capital $\mathcal{C}$ and illiquid assets $\psi$ with production cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ for dates $i\in\{0,1,\ldots,T\}$. The value $A^{\prime}_{i}\geq 0$ of the assets and the value $L_{i}\geq 0$ of the liabilities are defined by (with $\tilde{Z}_{t}^{\phi+\theta}\equiv\tilde{Z}_{t}^{\phi}+\tilde{Z}_{t}^{\theta}$):
$A^{\prime}_{i}=A^{\prime}_{i}(\mathcal{L},\phi,\psi,\mathcal{C})$ $=$ $\phi_{i}\cdot S_{i}+\tilde{Z}_{i}^{\phi+\psi+\mathcal{L}}$ (12)
$L_{i}=L_{i}(\mathcal{L},\phi,\psi,\mathcal{C})$ $=$ $X_{i}^{\mathcal{L}}+\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ (13)
@@
The fulfillment condition is a condition on the set $M_{i}=\{A^{\prime}_{i}\geq L_{i}\}\in\mathcal{F}_{i}$ ($i\in\{1,\ldots,T\}$) conditional on $\mathcal{F}_{i-1}$ with the property that if the set $M_{i}$ satisfies the condition, then so does any larger set $M\supseteq M_{i}$ in $\mathcal{F}_{i}$. The almost sure fulfillment condition corresponding to perfect production is $\mathbb{P}[M_{i}\mid\mathcal{F}_{i-1}]=1$.
For example (see Section 6), the fulfillment condition under Solvency II is $\mathbb{P}[M_{i}\mid\mathcal{F}_{i-1}]\geq 0.995$. Production strategies as in Definition 4.4 allow for capital $\mathcal{C}$ to be used in the production. As a simplification, we consider capital as one-year investments, with assets with market price $C_{i}\geq 0$ raised from investors at date $i$ and non-negative stochastic payoff $C_{i+1}^{\prime}$ to investors at date $i+1$ with which the capital investment arrangement is discharged. The payoff $C_{i+1}^{\prime}=(A^{\prime}_{i+1}-L_{i+1})_{+}$ as in Definition 4.3 reflects that claims of capital investors are subordinated to all other obligations and non-negative due to “limited liability”, and is based on the assumption that the relevant valuation standard corresponds to Definition 4.1. The financiability condition in Definition 4.3 below specifies when a stochastic payoff is sufficient for the capital investors. Capital investments can be seen as financial instruments in addition to tradables, and the definition below extends consistency for tradables in the sense of Definition 3.2 to include capital investments. This is in particular used in Section 5.1. It also introduces “neutrality to tradables”, which is in particular used in Section 5.3.
Definition 4.3 (Financiability condition, consistent with and neutral to tradables).
The financiability condition expresses, for $i\in\{0,1\ldots,T-1\}$, when the stochastic return from the capital investment
@@
is sufficient for the capital investors, with the following properties and definitions:
@@
    If $C_{i}\rightarrow C_{i+1}^{\prime}$ satisfies the financiability condition, then so does $C_{i}^{*}\rightarrow C_{i+1}^{*\prime}\,$ for any $C_{i}^{*}\leq C_{i}$ and $C_{i+1}^{*\prime}\geq C_{i+1}^{\prime}$.
  3. (c)
    The financiability condition is positively homogeneous iff, whenever $C_{i}\rightarrow C_{i+1}^{\prime}$ satisfies the financiability condition, then so does $\lambda_{i}\cdot C_{i}\rightarrow\lambda_{i}\cdot C_{i+1}^{\prime}$ for $\mathcal{F}_{i}$-measurable $\lambda_{i}\geq 0$.
  4. (d)
@@
Note that defining the return to the capital investors by $(A^{\prime}_{i+1}-L_{i+1})_{+}$ with $A^{\prime}_{i+1}$ defined for the illiquid assets $\psi$ as in Definition 4.1 means that we assume that the illiquid assets are not provided to the capital investors but “remain with the liabilities” $\mathcal{L}$. Further, intuitively, consistency with the tradables means that a capital investor would not invest more than the market price for the payoff of a self-financing strategy of tradables. Neutrality to the tradables intuitively means that a capital investor would accept to additionally invest in the payoff of a self-financing strategy of tradables for their market price.
We now discuss Definition 4.1 of the value of the assets and liabilities used in the fulfillment condition and its link to failure. In principle, a company fails when it defaults on a payment due or is balance sheet insolvent, i.e. the value of its assets is less than the value of its liabilities. In the framework, we define failure through balance sheet insolvency only, which is evaluated only at dates $i\in\{0,1,\ldots,T\}$, and defined by $A^{\prime}_{i}<L_{i}$, assuming that Definition 4.1 is the relevant valuation standard for the insolvency laws. We investigate these limitations in the following. We denote the value of the assets by $A^{\prime}_{i}$ instead of $A_{i}$ to indicate that the assets are before paying out the cash flows $X_{i}^{\mathcal{L}}$ at date $i$ and before any potential capital raise. If there is no failure until date $i$, i.e. $A^{\prime}_{i}\geq L_{i}$, the production strategy $\phi$ according to Definition 4.4 can be continued and so $\phi$ produces the liabilities “in sufficiently many cases”. In case of failure at date $i$, i.e. $A^{\prime}_{i}<L_{i}$, the strategy $\phi$ cannot be continued even after raising capital, as shown by the second part of Remark 4.7. In this case, the strategy $\phi$ defaults on the payment $X_{i}^{\mathcal{L}}$, or cannot produce the future cash flows $X_{t}^{\mathcal{L}}$ ($t>i$) “in sufficiently many cases”, i.e. default occurs in the future “more often” than allowed by the fulfillment condition. [^3] But is is possible that there is no default in the actually realized development. Balance sheet insolvency defined by $A^{\prime}_{i}<L_{i}$ thus also accounts for default. In principle, there may be another production strategy $\theta$ with lower production cost such that $A^{\prime}_{i}\geq L_{i}$, so we should in principle define the liabilities using a production strategy with minimal production cost.
We evaluate $A^{\prime}_{i}\geq L_{i}$ only at dates $i\in\{0,1,\ldots,T\}$, which means that no default is allowed to occur at dates $t\in D$ with $i<t<i+1$, i.e. the liability cash flows at these dates must always be produced. This simplification is not necessarily realistic and in a sense too strict.
The value $A^{\prime}_{i}$ of the assets at date $i$ according to Definition 4.1 contains only the cash flows $\tilde{Z}_{i}^{\psi}$ at this date and no other value for the illiquid assets $\psi$. This reflects that the illiquid assets are by assumption “fully illiquid”, in particular cannot be sold. However, as we see in Section 5.3, the future cash flows $\tilde{Z}_{t}^{\psi}$ for $t>i$ are reflected on the liability side in reducing the production cost $\bar{v}_{i}^{\phi,\psi}(\mathcal{L})$ of the liabilities.
As we now sketch and set out further below, the production of the liabilities $\mathcal{L}$ “in sufficiently many cases” with a production strategy $\phi$ according to Definition 4.4 proceeds recursively forward in time over successive one-year periods. At date $i\in\{1,\ldots,T-1\}$:
  1. 1.
    At the end of the production from $i-1$ to $i$, we have assets with value $A^{\prime}_{i}$ and liabilities with value $L_{i}$. In case $A^{\prime}_{i}<L_{i}$, failure occurs, which we investigate further in Section 4.2.
  2. 2.
    In the “sufficiently many cases” in which $A^{\prime}_{i}\geq L_{i}$, the liability cash flows $X_{i}^{\mathcal{L}}$ are paid and then the payoff $C_{i}^{\prime}=A^{\prime}_{i}-L_{i}$ is provided to the capital investors. (The payoff may be zero but overall provides a sufficient return from $i-1$ to $i$ according to the financiability condition.) Following this, the value of the liabilities is $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ and equal to the value of the assets: $A^{\prime}_{i}-X_{i}^{\mathcal{L}}-C_{i}^{\prime}=L_{i}-X_{i}^{\mathcal{L}}=\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$.
  3. 3.
    Including in the case $i=0$, with the financiability condition from $i$ to $i+1$ satisfied, additional assets with value equal to the capital $C_{i}$ are raised.
  4. 4.
    The production can be continued from $i$ to $i+1$ (see also Remark 4.7). That is, for all $t\in D$ with $i<t<i+1$, the liability cash flows $X_{t}^{\mathcal{L}}$ are paid and, at $i+1$, we have $A^{\prime}_{i+1}\geq L_{i+1}$ “in sufficiently many cases” according to the fulfillment condition.
@@
Let dates $i_{min}<i_{max}$ in $\{0,1,\ldots,T\}$, liabilities $\mathcal{L}$ with $\mathcal{F}_{i_{max}}$-measurable terminal value $Y_{i_{max}}$, a fulfillment and a financiability condition, and a set $M\in\mathcal{F}_{i_{min}}$ be given. A non-negative production strategy $\phi$ for producing the liabilities $\mathcal{L}$ with terminal value $Y_{i_{max}}$ on $D\equiv D[i_{min},i_{max}]$ and $M$ using capital $\mathcal{C}$ and illiquid assets $\psi$ is a non-negative strategy $\phi=(\phi_{t})_{t\in D}\in\mathcal{R}^{+}$ together with capital amounts $\mathcal{C}=(C_{i})_{i=i_{min}}^{i_{max}-1}$ with $C_{i}=C_{i}(\phi,\psi)\geq 0$ and illiquid assets $\psi$ such that, on $M$, for any $i\in\{i_{min},\ldots,i_{max}-1\}$, whenever $A^{\prime}_{i}\geq L_{i}$, the following conditions are satisfied:

### v2

@@ 3.1 Non-negative production strategies
Set-up. We consider an insurance company that is subject to regulatory solvency capital requirements and insolvency laws. Roughly, the insolvency laws determine the circumstances under which the company fails and the consequences of failure, and the regulatory solvency capital requirements intend to make failure sufficiently unlikely. We link this to sufficient fulfillment of the insurance liabilities by production strategies: at specified dates, no failure is intended to be equivalent to sufficient fulfillment according to the fulfillment condition (Definition 3.5 below) that is encoded in the regulatory solvency capital requirements. We assume in line with typical rules that a company fails when it defaults on a payment that is due or is balance sheet insolvent, i.e. the value of its assets is less than the value of its liabilities. We additionally assume a financiability condition (Definition 3.6 below) that specifies when the return from a capital investment is sufficient for capital investors.
@@
Insurance contracts. Let $\mathcal{L}$ denote the insurance contracts that the company has written until date 00, which we also sometimes imprecisely identify with the insurance liabilities from these contracts. The insurance contracts are assumed to be extinguished at date $T$, with no remaining cash flows or value, and we assume that the company writes no additional insurance contracts in the period $[0,T]$. We represent $\mathcal{L}$ by contractually specified cash outflows $X_{t}^{\mathcal{L}}\geq 0$ for claims and benefits and other e.g. administrative costs, and actual cash inflows $\tilde{Z}_{t}^{\mathcal{L}}\geq 0$ for premiums, at dates $t\in D[0,T]$. Our objective is to value the insurance liabilities $\mathcal{L}$ as a whole by the cost $\bar{v}_{0}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ of production strategies $\phi$ that sufficiently produce them according to the fulfillment condition, using capital $\mathcal{C}=(C_{i})_{i=0}^{T-1}$ and in general also illiquid assets $\psi$. As a simplification, we do not explicitly model the exercise of options by policyholders to terminate the insurance contracts and instead assume that this is implicitly taken into account in the cash flows $X_{t}^{\mathcal{L}}$ and $\tilde{Z}_{t}^{\mathcal{L}}$.
Production of insurance contracts. In the framework, insurance liabilities are fulfilled by production strategies that provide sufficient fulfillment. (Non-negative) production strategies as defined in the central Definition 3.7 below are more complicated than tradable investment strategies according to Definition 2.1 with cash inflows $\tilde{Z}_{t}^{\mathcal{L}}$ and outflows $X_{t}^{\mathcal{L}}$ in particular due to sufficient fulfillment, capital, and fully illiquid assets. In the present section, we only consider strategies that contain only non-negative units of tradables and that do not cover what happens in case of failure. We additionally distinguish between strategies with non-negative and negative production cost, for reasons that we explain below.
In line with the one-year view of regulatory capital requirements such as Solvency II and the SST, the production proceeds over successive one-year periods from $i-1$ to $i$ for $i=1,2,\ldots,T$, with recapitalization assumed to be possible at each annual date $i-1$. For simplicity, we only allow for production strategies under which the insurance company (almost surely) does not default on the payments $X_{t}^{\mathcal{L}}$ at dates $t\in D[0,T]\setminus\{0,1,\ldots,T\}$ and assume that balance sheet insolvency is only assessed at dates $i\in\{1,2,\ldots,T\}$. Hence, by assumption, failure is only possible at the latter dates.
At a date $i-1$, given no failure up to and including this date, the insurance company in general has a (non-negative) tradable asset portfolio $\phi_{\gamma(i-1)}$, an illiquid asset portfolio $\psi$, insurance liabilities $\mathcal{L}$, and capital arrangements with capital $C_{i-1}\geq 0$. From date $i-1$ to $i$, the company follows the given production strategy, which we denote by $\phi$. For this strategy, seen from date $i-1$, i.e. conditional on $\mathcal{F}_{i-1}$, the fulfillment condition is satisfied at $i$ and the financiability condition holds for the capital arrangement. Following the strategy, at each successive date $t\in D[0,T]$ with $i-1<t<i$, we assume the following sequence of steps: (i) cash inflows $\tilde{Z}_{t}^{\phi+\psi+\mathcal{L}}$ from the tradables $\phi$, the illiquid assets $\psi$, and the insurance contracts $\mathcal{L}$; (ii) the insurance obligations $X_{t}^{\mathcal{L}}$ are paid by selling tradables in $\phi_{t}$ or with the cash inflows $\tilde{Z}_{t}^{\phi+\psi+\mathcal{L}}$; and (iii) a new non-negative tradable portfolio $\phi_{\gamma(t)}$ is set up by potentially further trading tradables such that the following equality in terms of values holds: [^6] Recall the illiquid assets $\psi$ can by assumption not be traded.
Non-negativity (almost surely) in (12) corresponds to the above assumption that there is no default on the payments $X_{t}^{\mathcal{L}}$ for $i-1<t<i$.
At date $i$, the following additional aspects need to be taken into account: the company may fail; it may not be possible to continue the production strategy; and the value $C_{i}^{\prime}$ of the capital investment is assessed and potentially additional capital is raised. We make the following simplifying assumptions (which may not hold in reality due to different valuation standards): failure is defined by $A^{\prime}_{i}<L_{i}$ with the values $A^{\prime}_{i}$ of the assets and $L_{i}$ of the insurance liabilities according to Definition 3.1 below, and the same valuation determines the value $C_{i}^{\prime}$ of the capital investment. The value of the capital investment is then given by $C_{i}^{\prime}=(A^{\prime}_{i}-L_{i})_{+}$ because the claims of capital investors are subordinated to all other obligations and non-negative. Non-negativity is a consequence of the limited liability of the capital investors, i.e. in case of failure $A^{\prime}_{i}<L_{i}$, the capital investors get nothing but are also not obliged to inject additional capital. Note that, under the assumptions, the capital investment cannot cause failure.
We assume the following sequence of steps at date $i$ and then explain why specific steps are possible:
  1. (1)
    cash inflows $\tilde{Z}_{i}^{\phi+\psi+\mathcal{L}}$;
  2. (2)
    failure is assessed by comparing $A^{\prime}_{i}$ to $L_{i}$. In case of no failure $A^{\prime}_{i}\geq L_{i}$,
  3. (3)
    the insurance liability cash outflows $X_{i}^{\mathcal{L}}$ are provided;
  4. (4)
    the compensation $C_{i}^{\prime}$ is provided to the capital investors;
  5. (5)
    capital $C_{i}$ in the form of tradables is raised; [^7] Steps (4) and (5) are presented as separate steps for convenience. and
  6. (6)
    the new tradable portfolio $\phi_{\gamma(i)}$ is derived by potentially further trading tradables, satisfying:
The portfolio $\phi_{\gamma(i)}$ is the starting point for the production strategy for the next one-year period $i$ to $i+1$, which needs to be such that, conditional on $\mathcal{F}_{i}$, the fulfillment condition is satisfied at $i+1$, and the financiability condition is satisfied for the investment $C_{i}\rightarrow C_{i+1}^{\prime}$. The latter is why the capital $C_{i}$ can be raised in step 5, provided that moving to $\phi_{\gamma(i)}$ in step 6 is then possible. We explain below (Lemma 3.3 following the definition of the values $A^{\prime}_{i}$ and $L_{i}$ in Definition 3.1) why it is possible to move from step 2 to steps 3 and 4 and then to step 6. The two equations (12) and (13) also show that, in case of no failure, production strategies are investment strategies $\phi$ in the sense of Definition 2.1, with cash inflows $\tilde{Z}_{i}^{\psi+\mathcal{L}}+C_{i}$ and outflows $X_{i}^{\mathcal{L}}+C^{\prime}_{i}$ at date $i$.
As we explain below, the above sequence of steps only works for strategies with non-negative production cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})\geq 0$. To include negative production cost, which may e.g. occur when the outstanding premium cash inflows are relatively large, we introduce the following two additional steps 2bis between steps 2 and 3, and 5bis between 5 and 6:
  1. (2bis)
    tradables with value equal to $(-\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L}))_{+}$ are borrowed by taking on a corresponding liability.
  2. (5bis)
    the liability is paid back by tradables with value $(-\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L}))_{+}$ from the capital $C_{i}$.
This obviously works if it is possible to take on and close out additional liabilities in this way. Without this assumption, we are restricted to production strategies with non-negative production cost, as we discuss in more detail in Section 3.2. In the following, we thus distinguish two situations:
  1. (A)
    the above assumption is made and production cost can also be negative and steps 2bis and 5bis are included; or
  2. (B)
    the production strategies are restricted to non-negative production cost and steps 2bis and 5bis are not included.
Production cost. We define non-negative production cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})\geq 0$ at date $i$ as the (minimum) value of non-negative asset portfolios of tradables at $i$ such that, if such a portfolio is available prior to step 5 above, then production with the given strategy $\phi$ is possible from date $i$. In the above sequence of steps, one would then raise capital $C_{i}$ in step 5 and convert the tradables in step 6 to the tradable portfolio $\phi_{\gamma(i)}$. It follows that the production cost is the (known) difference between the value $v_{i}(\phi)=\phi_{\gamma(i)}\cdot S_{i}$ and the capital $C_{i}$, i.e. $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})=\phi_{\gamma(i)}\cdot S_{i}-C_{i}$, provided that this difference is non-negative.
Note that, although the production uses the illiquid assets $\psi$ and they in general reduce production cost in terms of the cost of the tradables required (Section 4.3), they are not part of the production cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$, which is limited to the value of the required tradables $\phi$.
For negative production cost, the corresponding starting point at date $i$ is a tradable portfolio with negative value, i.e. to start with a liability and no tradable assets. Negative production cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})<0$ at date $i$ can thus be defined as the negative of the (maximum) value of non-negative tradable portfolios that, in the additional step 5bis above, can be paid out immediately after raising the capital $C_{i}$ in step 5 above to settle the corresponding liability and then convert the remaining tradables with value $C_{i}-(-\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L}))_{+}$ in step 6 to the tradable portfolio $\phi_{\gamma(i)}$. It again follows that the production cost is given by $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})=\phi_{\gamma(i)}\cdot S_{i}-C_{i}$. Step 2bis and the need for steps 2bis and 5bis become clearer in the proof of Proposition 3.3 below.
Balance sheet valuation. We define the values $A^{\prime}_{i}$ and $L_{i}$ in the following by considering the situation between steps 1 and 2 above, i.e. prior to the cash outflows $X_{i}^{\mathcal{L}}$ at date $i$; this allows accounting for failure caused by default. We then discuss the definition and show in Proposition 3.3 below that no failure as defined by $A^{\prime}_{i}\geq L_{i}$ is equivalent to no default on $X_{i}^{\mathcal{L}}$ and the ability to continue the production strategy past date $i$. It follows from this in particular that steps 3 and 6 above are possible.
@@
Let $i\in\{1,2,\ldots,T\}$ and $\phi$ be a non-negative production strategy (Definition 3.7) for the liabilities $\mathcal{L}$ using capital $\mathcal{C}$ and illiquid assets $\psi$ with production cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$. The value $A^{\prime}_{i}\geq 0$ of the assets and the value $L_{i}\geq 0$ of the liabilities are defined by (with $\tilde{Z}_{t}^{\phi+\theta}\equiv\tilde{Z}_{t}^{\phi}+\tilde{Z}_{t}^{\theta}$):
$A^{\prime}_{i}=A^{\prime}_{i}(\mathcal{L},\phi,\psi,\mathcal{C})$ $=$ $\phi_{i}\cdot S_{i}+\tilde{Z}_{i}^{\phi+\psi+\mathcal{L}}+(-\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L}))_{+}$ (14)
$L_{i}=L_{i}(\mathcal{L},\phi,\psi,\mathcal{C})$ $=$ $X_{i}^{\mathcal{L}}+(\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L}))_{+}$ (15)
Remark 3.2.
Definition 3.1 contains all assets and liabilities at date $i$ in step 2 above explicitly or implicitly. The insurance contracts $\mathcal{L}$ are valued by their production cost as a liability when production cost are non-negative and as an asset otherwise. The illiquid assets $\psi$ and the cash inflows from $\mathcal{L}$, both of which by assumption can neither be bought nor sold, are accounted explicitly with their current cash inflows $\tilde{Z}_{i}^{\psi+\mathcal{L}}$ and implicitly in the production cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ to the extent that the future cash inflows from $\psi$ and $\mathcal{L}$ are used to produce the future insurance liability cash outflows and thus reduce the production cost (see Section 4.3).
We now show how failure links to default and continuing the production strategy:
Proposition 3.3.
Failure at date $i\in\{1,2,\ldots T\}$ as defined by $A^{\prime}_{i}<L_{i}$ using Definition 3.1 is equivalent to default on the payment $X_{i}^{\mathcal{L}}$ or the inability to continue the production strategy past date $i$.
Proof.
We have $L_{i}\geq X_{i}^{\mathcal{L}}$ by definition. First assume no failure, $A^{\prime}_{i}\geq L_{i}$. In case of non-negative production cost, the value of the total tradable assets is $A^{\prime}_{i}=\phi_{i}\cdot S_{i}+\tilde{Z}_{i}^{\phi+\psi+\mathcal{L}}$, so they are sufficient to pay the cash outflows $X_{i}^{\mathcal{L}}$. In case of negative production cost, however, the assets with value $(-\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L}))_{+}>0$ are not tradables, and given only $A^{\prime}_{i}\geq L_{i}$, the value $\phi_{i}\cdot S_{i}+\tilde{Z}_{i}^{\phi+\psi+\mathcal{L}}$ of the tradable assets is not necessarily sufficiently large to pay $X_{i}^{\mathcal{L}}$. This is where step 2bis above is used and links to the above definition of negative production cost: with the additional tradables with value equal to $-\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})>0$ borrowed in step 2bis, the total value of the tradables is $A^{\prime}_{i}$, so the cash outflows $X_{i}^{\mathcal{L}}$ can be paid and there is no default. To show that the production strategy can be continued, note that the value $C_{i}^{\prime}$ of the capital investment considers all assets and liabilities potentially including the borrowed tradables and the corresponding liability but these cancel each other out and so $C_{i}^{\prime}=A^{\prime}_{i}-L_{i}$. Inserting the definitions of $A^{\prime}_{i}$ and $L_{i}$, rearranging terms to isolate $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ on the right-hand side, and adding $C_{i}$ to both sides, we get $\phi_{i}\cdot S_{i}+\tilde{Z}_{i}^{\phi+\psi+\mathcal{L}}-X_{i}^{\mathcal{L}}-C_{i}^{\prime}+C_{i}=\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})+C_{i}=\phi_{\gamma(i)}\cdot S_{i}$ as required.
Finally, assume failure $A^{\prime}_{i}<L_{i}$, we similarly get $\phi_{i}\cdot S_{i}+\tilde{Z}_{i}^{\phi+\psi+\mathcal{L}}<X_{i}^{\mathcal{L}}+\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ and $C_{i}^{\prime}=0$. Possibly, there is then default on $X_{i}^{\mathcal{L}}$, but in any case, we get from the two previous relationships that $\phi_{i}\cdot S_{i}+\tilde{Z}_{i}^{\phi+\psi+\mathcal{L}}-X_{i}^{\mathcal{L}}-C_{i}^{\prime}+C_{i}<\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})+C_{i}=\phi_{\gamma(i)}\cdot S_{i}$, so the production strategy cannot be continued past date $i$. ∎
Remark 3.4.
In our framework, failure at a date $i$ means default at this date or not sufficient fulfillment after the date, i.e. default can occur in the future “more often” than allowed by the fulfillment condition. Proposition 3.3 does not directly show this, because failure $A^{\prime}_{i}<L_{i}$ is defined for a specific production strategy $\phi$ and there may be other production strategies for the production after date $i$ with lower production cost. This can be overcome by changing the production strategy after date $i$ to a strategy with minimal production cost. Because finding such a minimum can be challenging, we allow in the following also for production strategies that are not minimal in this sense and may thus overestimate production cost or provide more capital compensation than would be needed. In line with this, fulfillment and financiability condition are defined below as minimum requirements.
Definitions. We now define fulfillment and financiability condition and non-negative production strategies. Because of Proposition 3.3, sufficient fulfillment is defined in Definition 3.5 as a condition on the set on which $A^{\prime}_{i}\geq L_{i}$. Financiability conditions (Definition 3.6) express when the capital investment $C_{i-1}\rightarrow C_{i}^{\prime}=(A^{\prime}_{i}-L_{i})_{+}$ is acceptable to the capital investors. From now on, for convenience, we write a one-year period from date $i$ to $i+1$ instead of from $i-1$ to $i$.
@@
A fulfillment condition is a condition on the set $M_{i+1}=\{A^{\prime}_{i+1}\geq L_{i+1}\}\in\mathcal{F}_{i+1}$ ($i\in\{0,1,\ldots,T-1\}$) conditional on $\mathcal{F}_{i}$ with the property that if the set $M_{i+1}$ satisfies the condition, then so does any larger set $M\supseteq M_{i+1}$ in $\mathcal{F}_{i+1}$. The full fulfillment condition corresponding to full production is $\mathbb{P}[M_{i+1}\mid\mathcal{F}_{i}]=1$.
For example (see Section 5), the fulfillment condition under Solvency II is $\mathbb{P}[M_{i+1}\mid\mathcal{F}_{i}]\geq 0.995$.
Definition 3.6 (Financiability condition, consistent with and neutral to the tradables).
A financiability condition is a condition, for $i\in\{0,1,\ldots,T-1\}$, on when the stochastic return from the capital investment
@@
is sufficient for the capital investors, that has the following properties:
@@
    If $C_{i}\rightarrow C_{i+1}^{\prime}$ satisfies the financiability condition, then so does $C_{i}^{*}\rightarrow C_{i+1}^{*\prime}\,$ for any $0\leq C_{i}^{*}\leq C_{i}$ and $C_{i+1}^{*\prime}\geq C_{i+1}^{\prime}$.
We define:
  1. (1)
    A financiability condition is positively homogeneous iff, whenever $C_{i}\rightarrow C_{i+1}^{\prime}$ satisfies the financiability condition, then so does $\lambda_{i}\cdot C_{i}\rightarrow\lambda_{i}\cdot C_{i+1}^{\prime}$ for $\mathcal{F}_{i}$-measurable $\lambda_{i}\geq 0$.
  2. (2)
@@
Definition 3.6 also extends consistency for tradables from Definition 2.3 to include capital investments. This is in particular used in Section 4.1. Intuitively, consistency with the tradables means that a capital investor would not invest more than the market price for the payoff of a self-financing strategy of tradables. Neutrality to the tradables, which is in particular used in Section 4.3, intuitively means that a capital investor would accept to additionally invest in the payoff of a self-financing strategy of tradables for their market price.
We now define non-negative production strategies.
@@
Let the following be given: dates $i_{min}<i_{max}$ in $\{0,1,\ldots,T\}$, insurance contracts $\mathcal{L}$ with $\mathcal{F}_{i_{max}}$-measurable terminal value $Y_{i_{max}}$, a fulfillment and a financiability condition, and a set $M\in\mathcal{F}_{i_{min}}$. A non-negative production strategy $\phi\equiv\phi(\mathcal{C},\psi,D,M)$ for producing the liabilities $\mathcal{L}$ on $D\equiv D[i_{min},i_{max}]$ and $M$ using capital $\mathcal{C}$ and illiquid assets $\psi$ is a non-negative strategy $\phi=(\phi_{t})_{t\in D}\in\mathcal{R}^{\geq 0}$ together with capital amounts $\mathcal{C}=(C_{i})_{i=i_{min}}^{i_{max}-1}$ with $C_{i}=C_{i}(\phi,\psi)\geq 0$ and illiquid assets $\psi$ such that, on $M$, for any $i\in\{i_{min},\ldots,i_{max}-1\}$, whenever $A^{\prime}_{i}\geq L_{i}$ according to Definition 3.1, the following conditions are satisfied:
@@
If there is no tradable for which portfolios containing negative units of the tradable are available for production with close out (Lemma 2.4), production strategies must have non-negative production cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})\geq 0$.